OpenStudy (anonymous):

please help, 1 equation

OpenStudy (anonymous):

solve the following equation -5 = s - 3

OpenStudy (anonymous):

using step equations

OpenStudy (anonymous):

-5 = s - 3 Add 3 on both sides, what would yo get? @trollface

OpenStudy (anonymous):

8? @dpasingh

OpenStudy (anonymous):

@trollface no

OpenStudy (anonymous):

??

HanAkoSolo (jamierox4ev3r):

-5=s-3 just add three to both sides what is -5+3? @trollface

OpenStudy (anonymous):

-2

HanAkoSolo (jamierox4ev3r):

that is your answer, since all you have to do is solve for the variable. yes that is correct @trollface :)

OpenStudy (anonymous):

-5 = s - 3 Adding 3 on boith sides we find -5+3= s-3+3 i.e. -2 =s i.e. S= -2 is the correct answer. @trollface

OpenStudy (anonymous):

Ohhh, okay, I understand now..

HanAkoSolo (jamierox4ev3r):

yep :)

OpenStudy (anonymous):

Thanks so much!